# A Survey on Mixture of Experts

### v2

@@ Abstract
Large language models (LLMs) have garnered unprecedented advancements across diverse fields, ranging from natural language processing to computer vision and beyond. The prowess of LLMs is underpinned by their substantial model size, extensive and diverse datasets, and the vast computational power harnessed during training, all of which contribute to the emergent abilities of LLMs (e.g., in-context learning) that are not present in small models. Within this context, the mixture of experts (MoE) has emerged as an effective method for substantially scaling up model capacity with minimal computation overhead, gaining significant attention from academia and industry. Despite its growing prevalence, there lacks a systematic and comprehensive review of the literature on MoE. This survey seeks to bridge that gap, serving as an essential resource for researchers delving into the intricacies of MoE. We first briefly introduce the structure of the MoE layer, followed by proposing a new taxonomy of MoE. Next, we overview the core designs for various MoE models including both algorithmic and systemic aspects, alongside collections of available open-source implementations, hyperparameter configurations and empirical evaluations. Furthermore, we delineate the multifaceted applications of MoE in practice, and outline some potential directions for future research. To facilitate ongoing updates and the sharing of cutting-edge advances in MoE research, we have established a resource repository accessible at https://github.com/withinmiaov/A-Survey-on-Mixture-of-Experts.
@@
In the current landscape of artificial general intelligence (AGI), the transformative impact of transformer-based large language models (LLMs) has permeated diverse fields such as natural language processing [1, 2, 3, 4, 5, 6, 7], computer vision [8, 9], and multimodality [10, 11, 12, 13]. Building upon the foundational transformer architecture, LLMs demonstrate extraordinary capabilities, which are attributed to their sheer size, the breadth of data they are trained on, and the significant computational resources invested in their development [14, 15, 16]. Recognizing a scaling law [14, 17] that underpins their evolution, it is imperative to identify and implement efficient methodologies for the sustainable scaling of LLMs.
@@
Despite the increasing popularity and application of MoE models in various domains, the literature has yet to see a survey that thoroughly examines and categorizes the advancements in this area. The most recent review of MoE we could find was presented in September 2022 [34], predating the pivotal “ChatGPT moment”, which omits the significant advancements that have recently emerged alongside the escalating academic and industrial interest in this domain. This gap in the literature not only hinders the progress of MoE research but also limits the dissemination of knowledge on this topic to a broader audience. Our survey aims to address this deficit by providing a clear and comprehensive overview of MoE with a novel taxonomy that segments recent progress into algorithm, system, and application.
@@ 2.2 Sparse MoE
While a dense mixture of experts typically yields higher prediction accuracy [8], it also incurs a significant increase in computational overhead. To address this, Shazeer et al. [27] introduced the sparsely-gated MoE layer, which is designed to activate only a selected subset of experts during each forward pass. This strategy achieves sparsity by computing a weighted sum of the outputs from only the top- experts, rather than aggregating the outputs from all the experts. The structure of the sparse MoE layer is illustrated in Figure 2(b). Building on the framework established by [27], Equation (2) can be modified to reflect the sparsely-gated mechanism as follows:
@@
Although the sparse gate substantially expands the model’s parameter space without a corresponding increase in computational cost, it can lead to a load balancing issue. Such an issue refers to the uneven distribution of workload across experts, with some being frequently utilized and others seldom or never engaged. To address this, each MoE layer incorporates an auxiliary loss function that promotes an even distribution of tokens across experts within each batch, as described in many studies [28, 36, 35, 29, 68, 69, 67]. To formulate this concept, consider a batch of queries , comprising tokens, and experts indexed from to . Following [28, 36], the auxiliary load balancing loss for the batch is defined as
@@ 4.1.1 Sparse
Auxiliary Loss for Token-Choice Gating. Token-choice gating algorithms frequently incorporate an auxiliary loss during training to promote equitable token distribution across experts. Table I shows prevalent auxiliary loss functions leveraged in the field. Shazeer et al. [27] quantify the importance of an expert in relation to a training batch via the batchwise sum of the gate values for that expert. They define an additional loss , added to the overall loss function for the model. This loss, which is equal to the square of the coefficient of variation of the set of importance values and multiplied by a hand-tuned scaling factor , encourages all experts to have equal importance. Although promotes balance in importance, it does not guarantee an even distribution of training examples among experts, which can lead to execution inefficiencies on distributed computing environments. To address this, they introduce a second loss to ensure balanced loads. Building on this foundation, GShard [28] define a new differentiable auxiliary loss using a differentiable approximation (the dot-product of mean gates and mean gating decisions per expert), as detailed in Section 2.2. Switch Transformers [36] and many other subsequent studies [35, 29, 31, 68] have embraced this design, and enhancements [69, 33, 67] have been made to cater to diverse requirements. Nevertheless, ST-MoE [37] identified limitations with , particularly at larger scales, leading to unreliable training outcomes. To mitigate this, it introduces the integration of router z-loss , improving training stability without quality degradation by penalizing large logits entering the gating network. Since this loss encourages absolute magnitude of values to be smaller, roundoff errors are reduced, which can be quite impactful for exponential functions such as gating. Additionally, Mod-Squad [71] posits the difficulty of training multi-task models under such an expert-balancing loss, which may inadvertently force experts to set parameters on conflicting tasks or hinder the potential synergies from parameter sharing across complementary tasks. Instead, it proposes to maximize the mutual information (MI) between experts and tasks to build task-expert alignment. Differently, ModuleFormer [73] proposes to maximize the Mutual Information between experts and tokens. Furthermore, DS-MoE [64] extends the application of , calibrating different weightings , in Mixture-of-Attention (MoA, as illustrated in Figure 5 (a)) and FFN MoE modules of different size models.
@@ 4.2.1 Network Types
| Yuan 2.0-M32 [84] (2024) | 3.7B/40B | 2/32 | 2048 | 8192 | 24 | 16 | 256 | 1 | SwiGLU | 0 |
@@ 4.3.2 Attention
Recent studies have demonstrated the effectiveness of extending MoE to the attention layer [82, 83, 73]. Additionally, there is a new line of research has focused on the fusion of MoPE with the attention mechanism to enhance the model’s efficiency and adaptability. For instance, MoELoRA [47] applies MoE to the attention mechanism in a resource-efficient manner by leveraging a PEFT method, adopting LoRA [174] to construct the experts. Specifically, MoELoRA sets multiple LoRA experts to the Query and Value matrices of the attention mechanism, and utilizes a gating network to activate the top- experts related to the specific tasks during both training and inference phases. To alleviate routing randomness, MoELoRA employs a contrastive learning loss to control the training of experts. The contrastive learning loss is designed to accentuate the differences in output distributions between experts, thereby encouraging them to capture diverse features relevant to the downstream tasks. MoELoRA offers a solution for flexibly combining various computational modules tailored to downstream tasks.

### v3

@@ Abstract
Large language models (LLMs) have garnered unprecedented advancements across diverse fields, ranging from natural language processing to computer vision and beyond. The prowess of LLMs is underpinned by their substantial model size, extensive and diverse datasets, and the vast computational power harnessed during training, all of which contribute to the emergent abilities of LLMs (e.g., in-context learning) that are not present in small models. Within this context, the mixture of experts (MoE) has emerged as an effective method for substantially scaling up model capacity with minimal computation overhead, gaining significant attention from academia and industry. Despite its growing prevalence, there lacks a systematic and comprehensive review of the literature on MoE. This survey seeks to bridge that gap, serving as an essential resource for researchers delving into the intricacies of MoE. We first briefly introduce the structure of the MoE layer, followed by proposing a new taxonomy of MoE. Next, we overview the core designs for various MoE models including both algorithmic and systemic aspects, alongside collections of available open-source implementations, hyperparameter configurations and empirical evaluations. Furthermore, we delineate the multifaceted applications of MoE in practice, and outline some potential directions for future research. To facilitate ongoing updates and the sharing of cutting-edge advances in MoE research, we have established a resource repository at https://github.com/withinmiaov/A-Survey-on-Mixture-of-Experts-in-LLMs.
@@
In the current landscape of artificial general intelligence (AGI), the transformative impact of transformer-based large language models (LLMs) has permeated diverse fields such as natural language processing [1, 2, 3, 4, 5], computer vision [6, 7], and multimodality [8, 9, 10]. Building upon the foundational transformer architecture, LLMs demonstrate extraordinary capabilities, which are attributed to their sheer size, the breadth of data they are trained on, and the significant computational resources invested in their development [11, 12, 13]. Recognizing a scaling law [11, 14] that underpins their evolution, it is imperative to identify and implement efficient methodologies for the sustainable scaling of LLMs.
@@
Despite the growing popularity and application of mixture of experts (MoE) models across various domains, comprehensive reviews that thoroughly examine and categorize advancements, particularly in the context of MoE in LLMs, remain scarce. Specifically, we identified two surveys preceding our work: the first, published in August 2012 [31], provides a comprehensive review of early studies on dense MoE, which significantly differ from the current mainstream focus on sparse MoE; the second, released in September 2022 [32], predates the major developments following the “ChatGPT moment” and, as a result, does not cover the substantial advancements and increased interest from academia and industry that have since emerged. This gap in the literature not only hinders the progress of MoE research but also limits the broader dissemination of knowledge on this topic. Our survey aims to address this deficit by providing a clear and comprehensive overview of MoE in LLMs, introducing a novel taxonomy that organizes recent progress into three categories: algorithm, system, and application.
@@ 2.2 Sparse MoE
While dense MoE typically yields higher prediction accuracy [6], it also incurs a significant increase in computational overhead. To address this, Shazeer et al. [24] introduced the sparsely-gated MoE layer, which is designed to activate only a selected subset of experts during each forward pass. This strategy achieves sparsity by computing a weighted sum of the outputs from only the top- experts, rather than aggregating the outputs from all the experts. The structure of the sparse MoE layer is illustrated in Figure 2(b). Building on the framework established by [24], Equation (2) can be modified to reflect the sparsely-gated mechanism as follows:
@@
Although the sparse gate substantially expands the model’s parameter space without a corresponding increase in computational cost, it can lead to a load balancing issue. Such an issue refers to the uneven distribution of workload across experts, with some being frequently utilized and others seldom engaged. To address this, each MoE layer incorporates an auxiliary loss function that promotes an even distribution of tokens across experts within each batch, as described in many studies [25, 34, 33, 26, 66, 67, 65]. To formulate this concept, consider a batch of queries , comprising tokens, and experts indexed from to . Following [25, 34], the auxiliary load balancing loss for the batch is defined as
@@ 4.1.1 Sparse
Auxiliary Loss for Token-Choice Gating. Token-choice gating algorithms frequently incorporate an auxiliary loss during training to promote equitable token distribution across experts. Table I shows prevalent auxiliary loss functions leveraged in the field. Shazeer et al. [24] quantify the importance of an expert in relation to a training batch via the batchwise sum of the gate values for that expert, defined as
Furthermore, they introduce an additional loss , which is defined as the square of the coefficient of variation of the set of importance values, and can be formulated as
This loss is multiplied by a manually adjusted scaling factor , and then integrated into the overall loss function for the model, encouraging all experts to have equal importance. Although promotes balance in importance, it does not guarantee an even distribution of training examples among experts, which can lead to execution inefficiencies in distributed computing environments. To address this, they introduce a second loss , which employs a smooth estimator of the number of examples assigned to each expert for a batch of inputs, thereby facilitating gradient backpropagation. Simplifying the above design, GShard [25] defines a new differentiable auxiliary loss using a differentiable approximation (the dot-product of mean gates and mean gating decisions per expert). This is equivalent to in Equations (6)–(8), as detailed in Section 2.2. Switch Transformers [34] and many other subsequent studies [33, 26, 28, 66] have embraced this design, and enhancements [67, 30, 65] have been made to cater to diverse requirements. Nevertheless, ST-MoE [35] identified limitations with , particularly at larger scales, leading to unreliable training outcomes. To enhance training stability without compromising quality, the -loss is integrated, defined as
which functions by penalizing large logits entering the gating network. Since this loss encourages absolute magnitude of values to be smaller, roundoff errors are reduced, which can be quite impactful for exponential functions such as gating. Additionally, Mod-Squad [69] posits the difficulty of training multi-task models under such an expert-balancing loss, which may inadvertently force experts to set parameters on conflicting tasks or hinder the potential synergies from parameter sharing across complementary tasks. Instead, it proposes to maximize the mutual information (MI) between experts and tasks to build task-expert alignment, which is accomplished through . Differently, ModuleFormer [71] proposes to maximize the Mutual Information between experts and tokens. Furthermore, DS-MoE [62] extends the application of , calibrating different weightings , in Mixture-of-Attention (MoA, as illustrated in Figure 5 (a)) and FFN MoE modules of different size models. Although existing research has introduced many different gating functions and auxiliary losses, the methods proposed by GShard [25] remain the predominant choice in industry [33, 26, 28, 67, 66].
@@ 4.2.1 Network Types
| Yuan 2.0-M32 [82] (2024) | 3.7B/40B | 2/32 | 2048 | 8192 | 24 | 16 | 256 | 1 | SwiGLU | 0 | |
| OLMoE [173] (2024) | 1.3B/6.9B | 8/64 | 2048 | 1024 | 16 | 16 | 128 | 1 | SwiGLU | 0 | |
| DeepSeek-V3 [174] (2024) | 37B/671B | 9/257 | 7168 | 18432 | 2048 | 61 | 128 | 128 | 1, except first 3 layers | SwiGLU | 1 |
@@ 4.2.2 Hyperparameters
| OLMoE-0924-Base | 2024.9 | Ai2 | 1.3B | 6.9B | 54.1 | 3.0 (8-shot) | - | 22.4 | https://huggingface.co/allenai/OLMoE-1B-7B-0924 |
| DeepSeek-V3-Base | 2024.12 | DeepSeek | 37B | 671B | 87.1 | 89.3 (8-shot) | 61.6 | 65.2 | https://huggingface.co/deepseek-ai/DeepSeek-V3-Base |
@@ 4.3.2 Attention
Recent studies have demonstrated the effectiveness of extending MoE to the attention layer [80, 81, 71]. Additionally, there is a new line of research has focused on the fusion of MoPE with the attention mechanism to enhance the model’s efficiency and adaptability. For instance, MoELoRA [45] applies MoE to the attention mechanism in a resource-efficient manner by leveraging LoRA [172] to construct the experts. Specifically, MoELoRA sets multiple LoRA experts to the Query and Value matrices of the attention mechanism, and utilizes a gating network to activate the top- experts related to the specific tasks during both training and inference phases. To alleviate routing randomness, MoELoRA employs a contrastive learning loss to control the training of experts. The contrastive learning loss is designed to accentuate the differences in output distributions between experts, thereby encouraging them to capture diverse features relevant to the downstream tasks. MoELoRA offers a solution for flexibly combining various computational modules tailored to downstream tasks.